# Crude Approximation of Directed Minimum Cut and Arborescence Packing in Almost Linear Time

### v1

@@ 1.1 Our Results
Before stating our result on arborescence packing, we first define congestion. Let ${\cal T}=\{T_{1},\dots,T_{k^{\prime}}\}$ be a set of $s$-rooted arborescences. The congestion $\cal T$} is the maximum number of times any edge may appear in $T_{i}\in{\cal T}$, i.e., $\max_{e\in E}|\{T_{i}\in{\cal T}\mid e\in T_{i}\}|$.
In this work, we design an algorithm that achieves $n^{o(1)}$ congestion in almost-linear time. [^1] We note that our proof of Theorem 1.2 (see Theorem 5.2 for a detailed version) assumes an almost-linear-time directed expander routing subroutine (see Theorem 5.1), which is commonly believed, but we are not aware of a formal proof in the existing literature. We will work on a full proof in the next version.
@@
Independently, Quanrud [quanrud2025approximatingdirectedconnectivityalmostlinear] very recently showed a $(1+\epsilon)$-approximation algorithm for $s$-rooted connectivity using $m^{1+o(1)}$ time. His approximation of $(1+\epsilon)$ is stronger than our polylogarithmic approximation factor in Theorem 1.1. It is unclear if [quanrud2025approximatingdirectedconnectivityalmostlinear] implies any result in arborescence packing similar to Theorem 1.2.
@@ 4 Approximate rooted min-cut
3:  Compute mincut from $E^{-}(C)$ to random $v\leftarrow\text{endpoints of }E_{i}[C]$
4:  Repeat $\Theta(\log n)$ times
5:end for
6:return $\min(\rho(C_{v}))$ for all computed $C_{v}$
@@ 5.2.1 Invariants at Level $i$
    For each vertex $v\in V(G)\setminus\{s\}$, there is $|\Gamma_{i}(v)|\leq\Delta_{i}(v)\cdot(i+1)$.
@@
    For each edge $e\in E(G)$, there is $|\Gamma_{i}(e)|\leq 5i^{2}\cdot(\kappa_{\mathrm{route}}/\phi)$.
@@
We start by setting $\Gamma(e)=\Gamma_{L}(e)$ for each edge $e\in E(G)$. For each vertex $v\in V(G)\setminus\{s\}$, observe that $K_{L}(v)$ is exactly the inter-SCC incoming edges of $v$. We distribute the colors $\Gamma_{L}(v)$ over the edges in $K_{L}(v)$ such that each edge receives at most $L+1$ colors. Here, an edge $e\in K_{L}(v)$ receives a subset of colors $Y^{\prime}\subseteq\Gamma_{L}(v)$ means that we add $Y^{\prime}$ into $\Gamma_{i}(e)$. This finishes the computation of the final edge-coloring $\{\Gamma(e)\}$.
@@ 5.2.2 The Algorithm at Level $i$
We make some remarks on this partition. The first set $E_{X}$ is simply the level-$i$ critical incoming edges of $v$, which, by definition, consist of $v$’s incoming edges in $E_{>i}$ and $v$’s incoming edges whose tail-endpoints fall outside $C_{i}$. The second set $E_{Y}$ consists of $v$’s incoming edges whose tail-endpoints fall inside $C_{i}\setminus C_{i-1,v}$, but excludes those inside $E_{>i}$. By definition, $E_{X}$ and $E_{Y}$ are clearly disjoint. By the definition of $K_{i-1}(v)$, the remaining edges $K_{i-1}(v)\setminus(E_{X}\cup E_{Y})$ are all inside $E_{i}^{-}(v)$. Therefore, these three sets $E_{X},E_{Y},E_{Z}$ form a partition of $K_{i-1}(v)$.
@@
The remaining colors in $Z$ will be handled in the following steps, and we use $Z_{v}$ refer to the $Z$ for this vertex $z$.
@@
  • Suppose $\gamma\in Y_{u}$. In Step 1, $\gamma$ is assigned to an incoming edge $(v^{\prime},u)$ of $u$ such that $v^{\prime}\in C_{i}\setminus C_{i-1,v^{\prime}}$. This time, we repeat the above argument for the vertex $v^{\prime}$ and the color $\gamma$. Note that we make progress each time, as $C_{i-1,v^{\prime}}$ precedes $C_{i-1,v}$ in the order $\pi_{i}$, so finally we will reach the above two cases which find a $\gamma$-colored vertex in $C_{i}$ that can reach $v$ using only $\gamma$-colored edges.

### v2

# Approximating Directed Minimum Cut and Arborescence Packing via Directed Expander Hierarchies

@@ 1.1 Our Results
Before stating our next result, we first define the congestion of a set ${\cal T}=\{T_{1},\dots,T_{k^{\prime}}\}$ of $s$-rooted arborescences as the maximum number of times any edge may appear in $T_{i}\in{\cal T}$, i.e., $\max_{e\in E}|\{T_{i}\in{\cal T}\mid e\in T_{i}\}|$. Our second algorithm achieves $n^{o(1)}$ congestion in almost-linear time. [^1] We note that our proof of Theorem 1.2 (see Theorem 5.2 for a detailed version) assumes an almost-linear-time directed expander routing subroutine (see Theorem 5.1), which is a commonly believed, but we are not aware of a formal proof in the existing literature. We will work on a full proof in the next version.
@@
Independent of our work, Quanrud [quanrud2025approximatingdirectedconnectivityalmostlinear] and Mosenzon [mosenzon2025almost] have very recently obtained a $(1+\epsilon)$-approximation algorithm for $s$-rooted connectivity running in $m^{1+o(1)}/\epsilon$ time. This $(1+\epsilon)$ guarantee is stronger than our polylogarithmic approximation factor in Theorem 1.1. It remains unclear whether the techniques in [quanrud2025approximatingdirectedconnectivityalmostlinear, mosenzon2025almost] yield any analogous result for arborescence packing, as in Theorem 1.2.
@@ 4 Approximate rooted min-cut
3:  Compute mincut $\rho(C_{v})$ from $E^{-}(C)$ to random $v\leftarrow\text{endpoints of }E_{i}[C]$
4:  (When $i=0$, $C=\{v\}$ is a singleton and we get $\rho(C_{v})=\rho(v)$.)
5:  Repeat $\Theta(\log n)$ times
6:end for
7:return $\min(\rho(C_{v}))$ for all computed $C_{v}$
@@ 5.2.1 Invariants at Level $i$
    For each vertex $v\in V(G)\setminus\{s\}$, $|\Gamma_{i}(v)|\leq\Delta_{i}(v)\cdot(i+1)$.
@@
    For each edge $e\in E(G)$, $|\Gamma_{i}(e)|\leq 5i^{2}\cdot(\kappa_{\mathrm{route}}/\phi)$.
@@
We start by setting $\Gamma(e)=\Gamma_{L}(e)$ for each edge $e\in E(G)$. For each vertex $v\in V(G)\setminus\{s\}$, observe that $K_{L}(v)$ is exactly the inter-SCC incoming edges of $v$. We distribute the colors $\Gamma_{L}(v)$ over the edges in $K_{L}(v)$ such that each edge receives at most $L+1$ colors. Here, an edge $e\in K_{L}(v)$ receives a subset of colors $Y^{\prime}\subseteq\Gamma_{L}(v)$ means that we add $Y^{\prime}$ into $\Gamma(e)$. This finishes the computation of the final edge-coloring $\{\Gamma(e)\}$.
@@ 5.2.2 The Algorithm at Level $i$
We make some remarks on this partition. The first set $E_{X}$ is simply the level-$i$ critical incoming edges of $v$, which, by definition, consist of $v$’s incoming edges in $E_{>i}$ and $v$’s incoming edges whose tail-endpoints fall outside $C_{i}$. The second set $E_{Y}$ consists of $v$’s incoming edges whose tail-endpoints fall inside $C_{i}\setminus C_{i-1,v}$, but excludes those inside $E_{>i-1}$. By definition, $E_{X}$ and $E_{Y}$ are clearly disjoint. By the definition of $K_{i-1}(v)$, the remaining edges $K_{i-1}(v)\setminus(E_{X}\cup E_{Y})$ are all inside $E_{i}^{-}(v)$. Therefore, these three sets $E_{X},E_{Y},E_{Z}$ form a partition of $K_{i-1}(v)$.
@@
The remaining colors in $Z$ will be handled in the following steps, and we use $Z_{v}$ refer to the $Z$ for this vertex $v$.
@@
  • Suppose $\gamma\in Y_{u}$. In Step 1, $\gamma$ is assigned to an incoming edge $(v^{\prime},u)$ of $u$ such that $v^{\prime}\in C_{i}\setminus C_{i-1,v^{\prime}}$, and this edge $(v^{\prime},u)$ does not belong to $E_{>i-1}$. This time, we repeat the above argument for the vertex $v^{\prime}$ and the color $\gamma$. Note that we make progress each time, as $C_{i-1,v^{\prime}}$ precedes $C_{i-1,v}$ in the order $\pi_{i}$, so finally we will reach the above two cases which find a $\gamma$-colored vertex in $C_{i}$ that can reach $v$ using only $\gamma$-colored edges.